# Two-sample comparison through additive tree models for density ratios

### v2

@@ A.2 Derivation of the regularized prior for the node parameters
Hence, when $\gamma_{k}(A)=e^{\beta_{k}(A)}$ follows $\mathrm{InverseGauss}(1,K\lambda_{0})$, the leaf node parameter $\beta_{k}(A)$ approximately follows
@@
This observation implies that, under the prior
@@ B.1 MCMC algorithm
Updating the node parameters $\beta_{k}$ and the temperature $\tau$ is straightforward due to their conjugate structures. The algorithm to update $T_{k}$ with the Metropolis-Hastings algorithm is detailed in the next section.
@@ B.2 Updating trees
In this section, we consider updating the $k$th tree denoted by $T_{k}$ given the other parameter values using the proposals originally introduced for the Bayesian CART model (Chipman et al., 1998), and the details are obtained by following Kapelner and Bleich (2016).
@@
One of the moves is selected stochastically according to the fixed probabilities $p(\mathrm{GROW})$, $p(\mathrm{PRUNE})$, and $p(\mathrm{CHANGE})$, and they are set to 1/3 in our numerical experiments.
@@
$\alpha=\underbrace{\frac{p(T^{*}_{k})}{p(T_{k})}}_{\text{Prior}}\underbrace{\frac{p(T^{*}_{k}\to T_{k})}{p(T_{k}\to T^{*}_{k})}}_{\text{Transition}}\underbrace{\frac{p(\mathrm{data}\mid T_{k},p(\{w_{-k}(x_{i})\}^{n}_{i=1})}{p(\mathrm{data}\mid T^{*}_{k},p(\{w_{-k}(x_{i})\}^{n}_{i=1})}}_{\text{Loss-based likelihood}}$
@@
$\int p(\beta_{k})L_{n,\tau}(w)d\beta_{k}\propto\prod_{A\in\mathcal{L}(T_{k})}L_{k,A},$
where $p(\beta_{k})$ is the joint prior density of the leaf node parameters, $\mathcal{L}(T_{k})$ is the collection of the leaf nodes, and
$L_{A}=\sqrt{\frac{\lambda_{k}(A)}{\lambda^{\prime}_{k}(A)}}\exp\left(\frac{\lambda_{k}(A)}{\mu_{k}(A)}-\frac{\lambda^{\prime}_{k}(A)}{\mu^{\prime}_{k}(A)}\right).$
@@ C.3 Notes on the implementation of the density ratio estimators
For AdaBoost, we used the ada function from the ada package (Culp et al., 2016), and the optimal number of trees is selected via the $5$-fold cross-validation based on the classification error reported by the function. We found that the selected number tends to be close to zero when the two-sample difference is small (such as the local shift scenario in the 2D experiments), so we set the minimum number of trees to 10 manually.
@@
For the KLIEP and uLSIF, we use the functions KLIEP and uLSIF from the R package densratio (Makiyama, 2019) with the default parameter values. It should be noted that we detected an error in the source code for the KLIEP function 111 The estimated density ratios are not correctly computed on observations in a test data set in the original package. , so we used a fixed version that is available at https://github.com/nawaya040/densratio.
Appendix D Additional figures
@@
We provide additional visualizations corresponding to the settings described in Section 4.1. Figure 8 shows the estimated log-density ratios for the first (global shift) and third (local dispersion difference) scenarios under the balanced sample size setting, and Figure 9 shows the estimated ratios for the second (local shift) scenario under the unbalanced sample size configuration.
@@
Figure 8: The estimated log-density ratios obtained using the algorithms considered in the two-dimensional examples for the global shift scenario and the local dispersion scenario ($n_{0}=n_{1}=5000$). For the Bayesian additive model, the 2.5% quantiles and 97.5% quantiles of the log-density ratios evaluated point-wise are also displayed. The difference between FS and GB is minimal, so we present only the results for GB.
@@
Figure 9: The estimated log-density ratios obtained with the algorithms considered in the two-dimensional examples for the global shift scenario and the local dispersion scenario ($n_{0}=9000$ and $n_{1}=1000$). For the Bayesian additive model, the 2.5% quantiles and 97.5% quantiles of the log-density ratios evaluated point-wise are also displayed. The difference between FS and GB is minimal, so we present only the results for GB.

### v3

@@ A.2 Derivation of the regularized prior for the node parameters
Hence, when $\gamma_{k}(A)=\exp(\beta_{k}(A))$ follows $\mathrm{InverseGauss}(1,K\lambda_{0})$, the leaf node parameter $\beta_{k}(A)$ approximately follows
@@
This observation implies that, under the alternative prior
@@ B.1 MCMC algorithm
Updating the node parameters $\beta_{k}$ and the temperature $\tau$ is straightforward due to their conjugacy. The algorithm to update $T_{k}$ with the Metropolis-Hastings algorithm is detailed in the next section.
@@ B.2 Updating trees
In this section, we consider updating the $k$th tree denoted by $T_{k}$ given the other parameter values using the proposals originally introduced for the Bayesian CART model (Chipman et al., 1998), and the details are obtained by following Chipman et al. (2010); Kapelner and Bleich (2016).
@@
One of the moves is selected stochastically according to the fixed probabilities $p(\mathrm{GROW})$, $p(\mathrm{PRUNE})$, and $p(\mathrm{CHANGE})$, and they are all set to 1/3 in our numerical experiments.
@@
$\alpha=\underbrace{\frac{p(T^{*}_{k})}{p(T_{k})}}_{\text{Prior}}\underbrace{\frac{p(T^{*}_{k}\to T_{k})}{p(T_{k}\to T^{*}_{k})}}_{\text{Transition}}\underbrace{\frac{p(\mathrm{data}\mid T_{k},\{w_{-k}(x_{i})\}^{n}_{i=1})}{p(\mathrm{data}\mid T^{*}_{k},\{w_{-k}(x_{i})\}^{n}_{i=1})}}_{\text{Loss-based likelihood}}$
@@
$\int p(\beta_{k})L_{n,\tau}(w)\,d\beta_{k}\propto\prod_{A\in T_{k}}L_{k,A},$
where $p(\beta_{k})$ is the joint prior density of the leaf node parameters, and
$L_{k,A}=\sqrt{\frac{\lambda_{k}(A)}{\lambda^{\prime}_{k}(A)}}\exp\left(\frac{\lambda_{k}(A)}{\mu_{k}(A)}-\frac{\lambda^{\prime}_{k}(A)}{\mu^{\prime}_{k}(A)}\right).$
@@ C.3 Notes on the implementation of the density ratio estimators
For AdaBoost, we used the ada function from the ada package (Culp et al., 2016), and the optimal number of trees is selected via the $5$-fold cross-validation based on the classification error reported by the function. We found that the selected number tends to be extremely small when the two-sample difference is small (such as the local shift scenario in the 2D experiments), so we set the minimum number of trees to 10 manually. For the calibration method (Cranmer et al., 2015), we estimate the density functions of the classification probabilities for the two groups with the kernel density estimation implemented by the density function in R.
@@
For the KLIEP and uLSIF, we use the functions KLIEP and uLSIF from the R package densratio (Makiyama, 2019) with the default parameter values. It should be noted that we detected an error in the source code for the KLIEP function  [^1] The estimated density ratios are not correctly computed on observations in a test data set in the original package. , so we used a fixed version that is available at https://github.com/nawaya040/densratio .
Appendix D Additional numerical experiment results
D.1 Comparison with the density ratio trick in the one-dimensional scenarios
We perform a simple experiment to discuss how the bias can be introduced by the density ratio trick based on classification methods, if the prior information on the sample size is not effectively incorporated into the learning algorithms.
As in Section 3.4, we simulate data sets from the two Gaussian distributions, $N(0,1)$ and $N(1,2.25)$, and compare two algorithms: (i) the density ratio trick based on the Real AdaBoost with ada function from the ada package Culp et al. (2016) and (ii) the gradient boosting algorithm that optimizes the proposed balancing loss function. For both algorithms, the maximum tree depth and the size of ensemble are set to 2 and 100, respectively, and the learning rate is 0.01. We set the sample size $(n_{0},n_{1})$ to $(500,500)$, $(300,700)$, $(200,800)$, and $(100,900)$, changing the degree of unbalancedness, and the results are shown in Figure 8. We can see that the estimated log-odds ratio from AdaBoost is overall shrunk toward 0, which is the initial value in the algorithm (Culp et al., 2007), and it results in the positive bias in estimating the density ratios under the unbalanced settings. In contrast, for our proposed approach, even though the scale of the log-density ratio is underestimated due to the finite sample, the sign of the estimated overall agrees with that of the truth. The same tendency of the AdaBoost-based method introducing biases is observed in the two-dimensional cases, as illustrated in Figure 10.
Figure 8: A comparison of the density-ratio trick (DRT) approach based on the AdaBoost and the proposed gradient boosting (GB) under the 1D example with different sample sizes. The first column shows a comparison between the true posterior log-odds ratio and those estimated by the AdaBoost, and the second column visualizes the log-density ratios estimated with DRT and GB along with the truth. Their estimation errors are visualized in the third column. The estimated functions are the averages of the results based on 50 simulated data sets.
D.2 Additional results for the two-dimensional scenarios
@@
We provide additional visualizations corresponding to the settings described in Section 4.1. Figure 9 shows the estimated log-density ratios for the first (global shift) and third (local dispersion difference) scenarios under the balanced sample size setting, and Figure 10 shows the estimated ratios for the second (local shift) scenario under the unbalanced sample size configuration. For the unbalanced case, we note that the 95% credible intervals are above/below 0 for more than 10% of the observations, implying that the two-sample difference is significantly large in the corresponding regions, though this result is not clearly visible in the plots.
@@
Figure 9: The estimated log-density ratios obtained using the algorithms considered in the two-dimensional examples for the global shift scenario and the local dispersion scenario ($n_{0}=n_{1}=5000$). For the Bayesian additive model, the 2.5% quantiles and 97.5% quantiles of the log-density ratios evaluated pointwise are also displayed. The difference between FS and GB is minimal, so we present only the results for GB.
@@
Figure 10: The estimated log-density ratios obtained with the algorithms considered in the three scenarios with the unbalanced sample size ($n_{0}=9000$ and $n_{1}=1000$). For the Bayesian additive model, the 2.5% quantiles and 97.5% quantiles of the log-density ratios evaluated pointwise are also displayed. The difference between FS and GB is minimal, so we present only the results for GB.
We also show the calibration plots to evaluate the frequentist coverage rates of the credible intervals in Figure 11. We can observe only small over-coverage in the balanced scenarios, in which the data sets provide more information about the two-sample difference. In other words, we expect that the over-coverage and under-coverage under the unbalanced sample sizes are due to the structures of the differences that are difficult to accurately estimate with the small amount of information. For example, the diagonal shift in the global shift scenario is hard to describe with the axis-aligned tree partitions, and, for the local shift scenario, there is a sharp change in the density ratios in the local subset, but only a small number of observations from the second group (the smaller group) are observed there. Again, we note that such differences are accurately estimated under the balanced sample sizes in which richer information about the difference is provided from the data.
Figure 11: The calibration plots for the two-dimensional examples, in which the posterior uncertainty and the nominal coverage rates (namely, the ratio of observations included in the pointwise credible intervals) are compared under different probabilities. Each gray curve corresponds to one of the 50 simulated data sets, and the black curve indicates their averages.
D.3 Alternative simulation scenarios with 20-dimensional structures
In this section, we modify the simulation scenarios for the 20-dimensional cases so that the data structure becomes similar to the real microbiome data set analyzed in Section 5. To this end, we employ the following procedure:
  1. 1.
    We simulate a data set under the location shift scenario or the dispersion difference scenario.
  2. 2.
    We transform their marginal distributions by applying the CDFs of the Gaussian distributions  [^2] The parameters are selected so that the moments match those of the mixture component with the larger variances ($N(\mu_{2},\Sigma_{2})$). and then the inverse CDF of the beta distribution $\mathrm{Beta}(0.5,10)$.
As a result, the marginal distributions are right-skewed and have a large probability around 0, and the distributional structure is similar to the high-dimensional abundance data, as illustrated in Figure 12. The estimation results based on 50 simulated data sets are summarized in Table 3, and it shows that our proposed methods (BS, FS, and BAT) still estimate the two-sample difference more accurately.
Refer to caption Refer to caption
Figure 12: The first four marginal distributions of a simulated data set under the alternative 20-dimensional scenario based on “Location Shift”.
Table 3: A comparison of the MSEs and their standard errors in the alternative 20-dimensional scenarios.
Location Shift Dispersion
$n_{0}/(n_{0}+n_{1})$ 0.5 0.9 0.5 0.9
GB 0.077 0.129 0.163 0.236
(0.001) (0.002) (0.004) (0.005)
FS 0.080 0.138 0.171 0.248
(0.001) (0.002) (0.005) (0.006)
BAT 0.062 0.107 0.179 0.361
(0.001) (0.002) (0.004) (0.008)
DRT (AdaBoost) 0.243 3.651 0.149 2.940
(0.030) (0.152) (0.004) (0.163)
CDC (AdaBoost) 0.133 0.495 0.469 0.896
(0.008) (0.043) (0.024) (0.047)
KLIEP 0.682 0.681 0.538 0.539
(0.003) (0.004) (0.001) (0.001)
uLSIF 19.278 11.592 0.538 0.539
(7.435) (5.744) (0.001) (0.001)
  • Notes: The methods are GB (gradient boosting), FS (forward stagewise), BAT (Bayesian additive trees), DRT and CDC (density-ratio trick and calibrated discriminative classifier based on AdaBoost), and the two kernel-based methods KLIEP and uLSIF.
Appendix E Additional real data analysis results
Figure 13 and 14 provide the additional real data analysis results on the credible interval comparison between different generative models. The credible bands for all samples are shown in the PCoA plots.
Train v.s. Generated: Posterior mean of DRE
Train v.s. Generated: Lower 2.5% MCMC quantile of DRE
Train v.s. Generated: Upper 97.5% MCMC quantile of DRE
Test v.s. Generated: Posterior mean of DRE
Test v.s. Generated: Lower 2.5% MCMC quantile of DRE
Test v.s. Generated: Upper 97.5% MCMC quantile of DRE